OpenStudy (anonymous):

Select the equation of a line that is perpendicular to the line on the graph and passes through the point (3, 2).

OpenStudy (anonymous):

y = 1 over 3 x + 3 y = -1 over 3x + 3 y = 3 x + 2 y = - 1 over 3 x + 2